Sales personnel for Skillings Distributors submit weekly reports listing the customer contacts made during the week. A sample of 55 weekly reports showed a sample mean of 17.5 customer contacts per week. The sample standard deviation was 5.5. Provide 90% and 95% confidence intervals for the population mean number of weekly customer contacts for the sales personnel.

90% Confidence interval, to 2 decimals:

( , )

95% Confidence interval, to 2 decimals:

( , )
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = 17.5

sample standard deviation = s = 5.5

sample size = n = 55

Degrees of freedom = df = n - 1 = 55 - 1 = 54

At 90% confidence level the t is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.1

\alpha / 2 = 0.05 / 2 = 0.05

t\alpha /2,df = t0.05,54 = 1.674

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 1.674 * (5.5 / \sqrt 55)

= 1.24

The 90% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

17.5 - 1.24 < \mu < 17.5 + 1.24

16.26 < \mu < 18.74

90% Confidence interval (16.26 , 18.74)

At 95% confidence level the t is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,54 = 2.005

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.005 * (5..5 / \sqrt 55)

= 1.49

The 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

17.5 - 1.49 < \mu < 17.5 + 1.49

16.01 < \mu < 18.99

95% Confidence interval (16.01 , 18.99)